A girl kicked the ball with initial velocity of 20m/s at an angle of 40degrees with thehorizontal
Calculate the maximum height reached
Calculate the time of flight
Calculate zthe horizontal displacement of the projectile
1
Expert's answer
2021-10-05T10:06:13-0400

Height:


"H=\frac{v^2\sin^2\theta}{2g}=8.43\text{ m}."

Time:


"T=\frac{2v\sin\theta}{g}=2.6\text{ s}."

Displacement:


"D=\frac{v^2\sin(2\theta)}{g}=40.2\text{ m}."